If something is $110 and its on sale for 30 percent off how much would that item be?
Answer by mananth(16949)

30% off
Let the tag price be x
The sale price is 70 when tag price is 100
70 --------------100
110---------------x
70/110= 100/x
x=100*110/70
x= $157.14